OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1

OpenStudy (anonymous):

subtract the two and you get 2x=2 x=1 plug this into the other two equations and get y=3 so the solution is (1,3)

OpenStudy (anonymous):

Multiply the top equation by 2. This yields -4x + 2y = 2. Subtract the second equation from this one to solve for y. Then plug in y to solve for x.

OpenStudy (anonymous):

subtract the 2 equns. : 2x = 2 x = 1 and y = 1 + 2(1) = 3

OpenStudy (anonymous):

y=1+2x -4x+1+2x=-1 -2x=-2 x=1 y=1+2x so y=3

OpenStudy (anonymous):

(1,3)

OpenStudy (anonymous):

thanks guys :)